### v1

# Who Pays for Fairness? Rethinking Recourse under Social Burden

Abstract
Machine learning based predictions are increasingly used in sensitive decision-making applications that directly affect our lives. This has led to extensive research into ensuring the fairness of classifiers. Beyond just fair classification, emerging legislation now mandates that when a classifier delivers a negative decision, it must also offer actionable steps an individual can take to reverse that outcome. This concept is known as algorithmic recourse. Nevertheless, many researchers have expressed concerns about the fairness guarantees within the recourse process itself. In this work, we provide a holistic theoretical characterization of unfairness in algorithmic recourse, formally linking fairness guarantees in recourse and classification, and highlighting limitations of the standard equal cost paradigm. We then introduce a novel fairness framework based on social burden, along with a practical algorithm (MISOB), broadly applicable under real-world conditions. Empirical results on real-world datasets show that MISOB reduces the social burden across all groups without compromising overall classifier accuracy.
1 Introduction
The ubiquity of automated decision-making systems has considerably reshaped the landscape of how individuals access crucial resources and are granted important social and economic opportunities, ranging from securing credit to receiving essential public services [9, 42, 40]. In recent years, growing concerns have emerged regarding the fairness of these systems, as well as their societal impact [20, 26, 44]. Notably, instances have been documented in which algorithmic decision support systems produced discriminatory outcomes in domains such as welfare distribution [25] and criminal justice [1]. Moreover, in high-stakes scenarios, it is imperative that whenever a model yields a negative decision, it also provides the affected individual with clear and actionable recourse [28]. Unfortunately, affected individuals usually lack transparency and effective means to contest undesired decisions (e.g., bail denial) or take meaningful steps to improve their situation [^1] For example, independent audits have shown fairness concerns and limited contestability options in nationwide-deployed systems, such as COMPAS for recidivism prediction in the United States [1] and VioGén for gender-based violence risk assessment in Spain [17].. For this reason, the “right to recourse” is slowly emerging as a legal requirement in several legislations, e.g., in the European GDPR and the Artificial Intelligence (AI) Act [18].
Pursuing a given recourse often demands considerable effort from the affected individual. As noted by Ustun et al. [47], the average effort to achieve recourse can vary significantly across sensitive groups, raising important concerns about fairness in recourse. This disparity may arise when (i) the recourse itself differs depending on group membership, or when (ii) the recourse is formally identical, but the real-world effort required to carry it out varies in practice. For example, consider a method that offers suggestions to individuals denied a loan. It might recommend ”Increase your monthly income by \$500” to someone from a privileged group, while advising ”Maintain stable employment for 24 more months” to someone from an underprivileged background. The first recommendation is often more viable for individuals with access to high-paying jobs or stable employment opportunities. In contrast, maintaining long-term employment can be substantially more challenging for individuals in precarious jobs or members of low-income and otherwise underprivileged communities. Further, consider an algorithm suggesting that all rejected students retake standardized tests. It may impose minimal burden on those who can afford test preparation, tutoring, and multiple attempts, while students coming from low social and economic backgrounds often face barriers, including limited resources and time constraints due to many socio-economic factors [45]. Thus, what appears to be group-aware or uniform recommendations can translate into very different levels of effort, reinforcing structural inequalities.
(a) The separate treatment of fairness for predictive and recourse elements of the pipeline.
(b) Holistic metrics expose large disparities.
Figure 1: Conventional metrics to measure fairness in recourse hide disparities. (a) Even in a population balanced by sensitive group membership and ground‑truth outcomes (e.g., true ability to repay a loan), treating prediction and recourse fairness separately, as in the majority of prior work, can mask disparities: if positive classifications are unevenly distributed, one group bears recourse burdens more frequently even when per‑instance recourse costs are equal; misclassification (e.g., false negatives) further amplifies this effect. (b) Conventional cost‑parity metrics suggest balanced recourse costs across groups $S\in\{0,1\}$ (left), but our metric (Equation 2) reveals substantial disparity once group differences in decision and error rates are taken into account (right). The algorithms and dataset (CAU-LIN) are from Von Kügelgen et al. [48].
In light of these concerns, it is paramount to develop metrics to reliably quantify and monitor the fairness of recourse recommendations, along with practical methods for improving these metrics. Existing research has primarily examined unfairness in recourse by analyzing the costs imposed on individuals receiving negative (i.e.,“reject”) classifications, either at the individual or group level [48, 15, 29, 4, 50]. However, we argue that fairness in recourse depends not only on what type of recourse is provided but also on who is more likely to receive the reject decision in the first place. As an illustration, consider two groups of people $S=0$ and $S=1$ having the same number of individuals per group. If one of the groups (e.g. $S=1$) has higher acceptance rate than the other group, many more individuals from $S=0$ receive a “reject” decision. Widely utilized fairness-in-recourse metric – equal expected cost (of negatively predicted individuals) – will not capture the inherent unfairness in the system as we can balance the cost of recourse per individual across two groups (e.g. $10/100=100/1000=0.1$). However, individuals in group $S=0$ are disproportionately classified as negative ($1000$ individuals c.f. $100$) and therefore, a higher number of instances from $S=0$ bear the cost of recourse.
We emphasize that the above illustrated problem is amplified when we consider the ground-truth labels, e.g., an individual’s latent (and unobserved) ability to repay a loan instead of just classifier’s acceptance outcomes. Individuals may be wrongly classified and burdened with corrective actions, despite the error stemming from the decision support system itself. This misallocation of responsibility is troubling, particularly as error rates often differ systematically across sensitive groups, a well-documented challenge in algorithmic fairness [34]. All the issues mentioned above stem from the fact that current popular fairness-in-recourse paradigms do not correctly capture fairness issues in the whole system, which includes both the classifier’s components and the recourse’s components.
This paper presents a unified framework for assessing fairness in algorithmic recourse across the full decision pipeline, from initial classification through recourse recommendation (see Figure 1(a) for an overview). Our framework yields formal connections between prediction fairness and recourse fairness, showing how disparities in classification performance can propagate to recourse recommendations, and vice versa. Further, by leveraging these insights, we introduce a lightweight training method, MISOB (MInimax SOcial Burden), that empirically improves fairness throughout the pipeline, without requiring access to sensitive attributes, in several real-world datasets. Lastly, we argue for moving beyond simple group-gap metrics and developing richer measures that more fully describe fairness guarantees.
Our contributions. Summarizing, we (i) theoretically charaterize the sources of unfairness in algorithmic recourse, establishing formal connections between fairness guarantees in recourse and classification (Section 3), (ii) provide evidence of the limitations of approaches based on the standard equal cost paradigm (Section 3.1), (iii) introduce a novel framework for fair recourse (Algorithm 1) based on the concept of social burden (Section 4), and (iv) assess its performance on real-world datasets (Section 5).
Table 1: Summary of the characteristics of existing works on fairness in recourse. We report the proposed fairness metric, whether it considers the classifier ($f(\mathbf{x})$) or the recourse method ($g_{f}(\mathbf{x})$), and the approach used to enhance fairness; specifically, whether it is agnostic to the classifier, the recourse method, or the sensitive attributes ($\mathcal{S}$). A “-” indicates works that propose only evaluation metrics without introducing any algorithm or methodology improving fairness. Our proposed metrics (Section 3) provide a holistic assessment of all components of the pipeline, whereas our approach (Section 4) improves them by remaining agnostic to the underlying components.
Fairness Metric Approach agnostic to
$f(\mathbf{x})$ $g_{f}(\mathbf{x})$ $Y$ $f(\mathbf{x})$ $g_{f}(\mathbf{x})$ $\mathcal{S}$
Gupta et al. [22]
Von Kügelgen et al. [48]
Raimondi et al. [43] - - -
Kuratomi et al. [30] - - -
Kavouras et al. [29] - - -
Bell et al. [4]
This Work (Section 3 and 4)
2 Preliminaries and Related Work
Notation and Problem Formulation.
Each decision subject (i.e., instance) is represented by the triplet $(\boldsymbol{x},\boldsymbol{s},y)$, where $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ represents the non-sensitive attributes, $\boldsymbol{s}\in\mathcal{S}\subseteq\mathbb{R}^{m}$ the sensitive demographic attributes (e.g., race, gender, age, marital status) and $y\in\mathcal{Y}$ the class label. For simplicity, we assume a binary classification task with $\mathcal{Y}=\{0,1\}$ (e.g., whether the decision subject is capable of repaying the loan). However, the proposed framework is general and can be extended to multiclass settings. We use $P$ to denote the probability of specific events, and $p$ to refer to probability distributions. The sensitive attribute vector $\boldsymbol{s}$ may be multidimensional, allowing individuals to be characterized by multiple protected attributes simultaneously. This enables the analysis of intersectional groups, e.g., “indigenous women under 30”, which capture more nuanced and realistic socio-economic contexts, often subject to distinct and compounded forms of disadvantage.
Further, let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a binary classifier whose goal is to assign a label in a set $\mathcal{Y}$ to instances in a set $\mathcal{X}$. Let $\delta((\boldsymbol{x},\boldsymbol{s}),\boldsymbol{x}^{\prime})$ denote the cost for an individual with features $(\boldsymbol{x},\boldsymbol{s})$ to change its non-sensitive profile to $\boldsymbol{x}^{\prime}$, e.g., $\ell_{1}$-norm [28]. Given an instance that received a negative classification (i.e., $f(\boldsymbol{x})=0$), we want to find the “closest” instance $\mathbf{x}^{\prime}$ achieving a positive classification by solving the following optimization problem:
$\mathop{\mathrm{argmin}}_{\mathbf{x^{\prime}}\in\mathcal{X}}\;\delta((\boldsymbol{x},\boldsymbol{s}),\mathbf{x}^{\prime})\quad\text{s.t.}\;f(\mathbf{x^{\prime}})\neq f(\mathbf{x})$ (1)
In the literature, $\mathbf{x}^{\prime}$ is also called counterfactual explanation [49]. In our paper, we assume the recommendation $\mathbf{x}^{\prime}=g_{f}(\boldsymbol{x})$ is generated by a recourse algorithm $g_{f}:\mathcal{X}\rightarrow\mathcal{X}$ tailored to the classifier $f$. In algorithmic recourse, a growing body of work has proposed diverse strategies $g_{f}(\boldsymbol{x})$ to offer recourse to individuals assigned unfavorable predictions. These methods vary in several key aspects, including the type of base classifier [47], the optimization technique [11, 36], and the incorporation of actionability or plausibility constraints [27, 10]. Our discussion here is clearly limited; thus, please refer to Karimi et al. [28] for a comprehensive survey on the topic. While providing recommendations is an important step toward accountability and transparency, significant fairness challenges remain, some stemming directly from the recourse process. For example, recourse can exacerbate social segregation [19], and its cost may vary with user preferences [10, 16] and sensitive attributes [50], potentially deepening existing disparities.
Measuring Algorithmic Fairness.
Research on algorithmic fairness has primarily examined disparities in model predictions across population subgroups, focusing on the fairness of predictions themselves. Fairness definitions typically fall into two categories: individual-level, which requires similar individuals to receive similar outcomes, and group-level, which evaluates whether performance metrics (e.g., accuracy, true positive rate (TPR)) are comparable across demographic groups. Group metrics are widely used due to their computational simplicity, ease of integration into machine-learning pipelines, and ability to reveal structural disparities. Two common examples are statistical parity [14], which measures differences in the probability of a positive prediction (i.e., acceptance rate (AR)) across groups regardless of true labels, and equality of opportunity [24], which measures such differences only among individuals in the positive class (i.e., TPR). However, these metrics focus solely on the prediction stage and overlook the downstream consequences of model decisions.
Measuring and Enhancing Fairness in Recourse.
Fairness concerns in algorithmic recourse prompted the search for specific fairness criteria [47, 22]. Table 1 summarizes the characteristics of the most relevant approaches that address fairness issues in recourse. Many works proposed both group- and individual-level fairness metrics in this direction [48, 29, 4] and extended the analysis to incorporate classifier predictions and standard predictive performance measures [30, 43]. Several papers that introduced fairness metrics for evaluating algorithmic recourse have also proposed methods to improve fairness according to these metrics [22, 48, 4]. For instance, Von Kügelgen et al. [48] proposed a causality-based method that requires access to a structural causal model (SCM). Gupta et al. [22] introduced an approach applicable only to recourse methods that map instances directly to the decision boundary, while Bell et al. [4] addressed fairness under resource constraints, a setting closer to a ranking rather than a classification problem.
Most fairness metrics proposed in this context focus exclusively on the recourse stage, considering only negatively classified instances and ignoring the classifier’s overall behavior, such as group-specific acceptance and rejection rates. Furthermore, they generally lack practical strategies for improving fairness. None of these studies adopt an intersectional perspective, and all assume access to sensitive attributes for all individuals at training time, an assumption rarely satisfied in real-world applications. Our work is conceptually aligned with Kuratomi et al. [30] and Raimondi et al. [43], but their formulation lacks a clear connection to existing fairness metrics in prediction. Further, these studies neither propose practical algorithms to enhance it nor address the importance of intersectional groups in evaluating recourse fairness.
3 Revisiting (Un)Fairness in Recourse
In this section, we outline three key limitations of common fairness metrics in pipelines with recourse: (1) they ignore the classifier’s decision behavior, (2) they omit ground truth information, and (3) they focus solely on metric equalization. We propose an alternative formulation that addresses these issues, offers a broader view of unfairness sources in recourse, and establishes explicit links between fairness in prediction and recourse. Proofs are provided in Appendix A.
3.1 The Impact of the Classifier’s Decisions
Let us exemplify the first issue with a simple example of a binary classification task. Consider a population $X$ with a binary target variable $Y\in\{0,1\}$ and two equally represented sensitive groups $S\in\{0,1\}$, such that $P(S=0)=P(S=1)$ and $P(Y=1\mid S=0)=P(Y=1\mid S=1)$. Let $f(x)$ be a classifier approximating the posterior $P(Y\mid X)$, and assume that under $f$ one group has a higher AR, i.e., $P(f(\boldsymbol{x})=1\mid S=1)>P(f(\boldsymbol{x})=1\mid S=0)$. After recourse is provided to negatively classified instances, let us assume that the expected cost of the required changes is equal across groups: $\mathbb{E}[\delta((\mathbf{x},\mathbf{s}),g_{f}(\mathbf{x}))\mid S=1]=\mathbb{E}[\delta((\mathbf{x},\mathbf{s}),g_{f}(\mathbf{x}))\mid S=0]$. According to standard fairness metrics, this situation would be considered fair. But is it truly fair? While the expected recourse cost is equal once offered, individuals in the group with the lower AR are more likely to require recourse in the first place. We argue that this disparity reflects an unfair situation, which existing popular metrics fail to capture. We now introduce an alternative fairness metric that evaluates recourse costs across the entire population, rather than only among negatively classified instances.
Definition 3.1.
Let $f$ be a classifier and $g_{f}$ the recourse algorithm. Let $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))$ be the cost associated to $\boldsymbol{x}$ of applying the recourse provided by $g_{f}$. Further, consider the expected recourse cost of the sensitive group $\boldsymbol{s}$ as $\mathbb{E}_{P(\boldsymbol{X}=\boldsymbol{x}|S=\boldsymbol{s})}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]$. Then, the expected cost in terms of the classifier’s AR i.e., $P(f(\boldsymbol{x})=1\mid S=\mathbf{s})$, can be defined as:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]\cdot(1-P(f(\boldsymbol{x})=1\mid S=\mathbf{s}))$ (2)
where the expectation is over $p(X\mid S=\boldsymbol{s},f(\boldsymbol{x})=0)$.
Namely, Equation 2 arises since instances receiving a positive prediction, $P(f(\boldsymbol{x})=1\mid S=\mathbf{s})$, do not modify their non-sensitive features, implying no changes in recourse cost, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$.
Figure 1(b) empirically proves the shortcomings of traditional fairness recourse metrics, that capture only the left-hand factor of Equation 2, potentially giving a misleading impression of fairness. Given the previous synthetic binary classification task, we computed the original metric from Von Kügelgen et al. [48] (left), averaging cost only over negatively predicted instances, and our revised formulation from Equation 2, which also accounts for the AR (right) of the base classifiers. When accounting for AR across sensitive groups, disparities in expected recourse costs become substantially larger. For instance, with svm, group S=1S=-1 has the lowest cost under the original metric but the highest under the revised one, as a significantly larger proportion of this group receives negative predictions. Therefore, to avoid masking underlying unfairness, we argue that the cost metric must account for acceptance rates to accurately capture true recourse costs across groups.
3.2 The Burden of False Negatives
In Equation 2, we measure the expected recourse cost while accounting for the classifier’s predictions, providing a more precise notion of fairness. However, it focuses exclusively on the model’s predictions $f(\mathbf{x})$, without accounting for the true class label. As a result, it fails to distinguish between individuals who were incorrectly denied a positive outcome (i.e., truly positive but misclassified) and those who were correctly assigned a negative outcome. In this section, we take a step further by focusing on minimizing the expected cost associated with positive instances that the classifier incorrectly predicts as negative. This next step naturally leads to the concept of social burden. [^2] Some works on algorithmic recourse (e.g., Kavouras et al. [29]) employ the term burden to refer to the average cost itself. However, in an earlier work by Milli et al. [35], burden is defined as the average cost incurred by instances that have a true positive class label. In this paper, we adopt the latter definition.
Social burden measures the expected cost incurred by individuals who are already qualified (true positives) but receive an unfavorable prediction by a classifier $f$ [35]. The social burden is zero if the classifier positively classifies all qualified individuals without requiring changes to their non-sensitive attributes. Conversely, the social burden is greater than zero when qualified individuals are assigned a negative classification and forced to alter their features to be recognized as valid. Formally, we can define social burden as:
Definition 3.2.
Let $f$ be the classifier, $g_{f}$ the recourse algorithm, and $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))$ the cost function. We denote the expected social burden for the sensitive group $\boldsymbol{s}$ as $\mathbb{E}_{P(\boldsymbol{X}=\boldsymbol{x}|S=\boldsymbol{s},Y=1)}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]$. Then, the expected cost in terms of the classifier’s true positive rate can be defined as:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}[\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))]\cdot(1-P(f(\boldsymbol{x})=1\mid S=\boldsymbol{s},Y=1))$ (3)
with the expectation over $p(X\mid S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)$.
Similarly to Equation 2, the instances obtaining a positive prediction do not count towards the overall burden.
Equation 3 underscores the importance of considering the predictive performance of the base classifier when evaluating algorithmic recourse, a factor often overlooked in the recourse literature but central to algorithmic fairness; especially, since numerous studies have shown that machine learning classifiers tend to exhibit unequal error rates (e.g., TPR) across sensitive groups [41]. When error rates differ across groups, the burden of recourse is also likely to be unevenly distributed, potentially amplifying existing disparities. However, it is important to note that satisfying equality of opportunity (i.e., parity in TPR) does not guarantee a small gap in social burden.
3.3 Gap-Based Metrics can Mask Unfairness
Unfairness in recourse can be quantified by measuring gaps in cost or social burden across sensitive groups (e.g, $\Delta\mathcal{C}_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|\mathcal{C}^{s}_{f,g}-\mathcal{C}^{s^{\prime}}_{f,g}|$ or $\Delta\mathcal{B}_{f,g}=\max_{s,s^{\prime}\in\mathcal{S}}|\mathcal{B}^{s}_{f,g}-\mathcal{B}^{s^{\prime}}_{f,g}|$). According to Equations 2 and 3, these disparities stem from three main sources. For cost gaps, the causes include: (a) disparities in AR, as captured by statistical parity [14]; (b) differences in the cost function, which may arise from unequal opportunities or socio-economic conditions; and (c) variations in the distribution $P(X\mid S=\boldsymbol{s},f(\boldsymbol{x})=0)$, reflecting potential population skews. For social burden gaps, they arise from: (a) disparities in true positive rates, captured by equality of opportunity [23], (b) unequal cost functions, and (c) skews in the distribution of qualified individuals denied recourse (i.e., $P(X\mid S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)$). Importantly, this illustrates that satisfying statistical parity (equality of opportunity) does not guarantee an equal distribution of recourse costs (social burden) across groups.
Although comparing the gap of statistical measures across groups is a common approach in the fairness literature, it can obscure meaningful underlying unfairness. In fairness for prediction, a well-known limitation of parity-based definitions is that parity can be achieved by degrading performance for the privileged group without improving outcomes for the unprivileged one [33, 13]. As a result, groups may exhibit similar expected values for a given metric, even when those values are far from optimal. In recourse, this issue is especially critical when evaluating social burden, which ideally should be as low as possible (preferably zero). A small gap in burden does not necessarily indicate fairness, as parity might be reached by increasing the burden for the privileged group rather than reducing it for the unprivileged one. However, in many real-world contexts, such as access to essential goods and services, sacrificing overall well-being is not acceptable. Therefore, fairness efforts should aim to reduce the burden for all groups, rather than simply equalizing it.
We propose drawing inspiration from approaches in the fairness-in-prediction literature that address this limitation and complement difference-based disparity measures with a Rawlsian (i.e., minimax) perspective [5]. For instance, in addition to evaluating the gap in social burden, we can examine the worst-group social burden defined as $\max_{\boldsymbol{s}\in\mathcal{S}}\mathcal{B}_{f,g}^{\boldsymbol{s}}$. This perspective’s benefits are twofold: (a) it helps identify the most vulnerable group, and (b) it serves as a base for the development of fairness-enhancing recourse frameworks that go beyond minimizing burden disparities, aiming instead to reduce the social burden for all groups.
Algorithm 1 MISOB: MInimax SOcial Burden
[…]
1: Pre-train $f^{(0)}$ using $\mathcal{D}$ // Warm-up Phase
2:for $t=1$ to $T$ do
3:  $\mathcal{Q}=\{\}$
4:  for $i=0$ to NN do
5:   $y^{i}\leftarrow f^{(t-1)}(\boldsymbol{x}^{i})$ // Predict label
6:   $b_{f,g}^{i}\leftarrow\delta(\boldsymbol{x}^{i},g_{f}(\boldsymbol{x}^{i}))\mathbb{I}\{y^{i}=1\}$ // Compute burden
7:   $\mathcal{Q}\leftarrow\mathcal{Q}\cup b_{f,g}^{i}$
8:  end for
9:  $f^{(t)}\leftarrow\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\phi(i,\mathcal{Q})\cdot\ell(f(\boldsymbol{x}^{i}),y^{i})$
10:end for
11:return  $f^{(T)}$
4 MISOB: MInimax SOcial Burden
In this section, we introduce a simple yet effective iterative training procedure, MISOB (MInimax SOcial Burden), that empirically reduces social burden (Equation 3) without degrading predictive performance and without requiring access to sensitive group attributes during training or inference. The method is outlined in Algorithm 1. Given a pre-trained classifier (line 1) [^3] Empirically, we observed that pretraining a base classifier to achieve reasonable accuracy prior to applying MISOB contributes to a more stable optimization process., the key idea is to incorporate social burden estimates directly into the training loop. At each iteration, the current classifier is used to generate predictions (line 5), from which the social burden for each instance is computed (line 6), and saved within $\mathcal{Q}$ (line 7). The model is then updated by minimizing a (weighted) loss function $\ell(f(\boldsymbol{x}^{i}),y^{i})$ (line 9), where weights are determined by the estimated burden $\phi(i,\mathcal{Q})$. Specifically, we propose assigning weights to instances $\boldsymbol{x}^{i}$ based on the following scheme:
$\phi(i,\mathcal{Q})=\begin{cases}1+CN\dfrac{b_{f,g_{f}}^{i}}{\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}}&\text{if }B>0\\ 1&\text{if }B=0\end{cases}$ (4)
where $N$ is the dataset (or batch) size, $B=\sum_{j\in|\mathcal{Q}|}b_{f,g_{f}}^{j}$ is the total burden with $b_{f,g}^{i}=\delta(\boldsymbol{x}^{i},g_{f}(\boldsymbol{x}^{i}))\cdot\allowbreak\mathbb{I}\{y^{i}=1\}$ being the burden of instance $i$, and $C$ is a hyperparameter controlling the influence of the burden term. Under this weighting scheme, each instance $\boldsymbol{x}^{i}$ is assigned a weight proportional to its contribution to the total burden across the dataset (or batch). In the degenerate case, where the total burden is zero, all instances are assigned equal weights. This reweighting scheme amplifies the influence of high-burden instances during training, encouraging the classifier to prioritize decisions that reduce unnecessary or costly adjustments imposed on individuals under the current decision boundary.
The proposed framework offers several advantages that enhance its practicality and broad applicability. [^4] The runtime of Algorithm 1 is determined by three factors: (i) the number of retraining steps $T$, (ii) the dataset size $N$, and (iii) the run-time complexity of the recourse method $O(K)$. In practice, datasets are typically much larger than the number of retraining iterations, and if a simple recourse method is used, we generally have $K<<T<<N$. As a result, the overall computational complexity is $O(N^{3})$. Efficiency can be further improved through batching, parallelization, or by approximating recourse computations. First, the method is agnostic to both the base classifier and the chosen recourse strategy, allowing seamless integration with a wide range of existing approaches. It also operates without access to sensitive attributes during training, avoiding legal, ethical, and practical challenges associated with collecting or using such data. Further, MISOB not only improves deployability but also naturally addresses intersectional fairness. Because group definitions are not required a-priori, intersectional subgroups can be introduced post-hoc at evaluation time, enabling fairness assessments across multiple dimensions without retraining. This flexibility makes the framework particularly well-suited for real-world applications where fairness concerns span diverse and evolving group identities.
5 Empirical Evaluation
In this section, we evaluate the effectiveness of MISOB in reducing the social burden across sensitive groups without compromising the overall predictive performance of the classifier. [^5] We will release the source code on GitHub under a permissive license. To this end, we experiment with various datasets, base classifiers and recourse methods, assessing how integrating burden information into the training process improves fairness guarantees in the algorithmic pipeline under study.
5.1 Experimental Setting
Datasets.
We consider the real-world dataset Adult [3] with sensitive groups defined by different characterizations of the sensitive attributes (race and gender), considering also intersectional groups. For results with other datasets, refer to Appendix C.
Base model.
For the base model, we consider a neural network (NN) with 2 hidden layers composed of 128 neurons each. We have chosen this base classifier because it can be merged with multiple different recourse methods proposed in the literature. Experiments with additional base classifiers can be found in Appendix C. We combine the NN with MISOB as well as with the method to enhance prediction fairness in terms of equality of opportunity (i.e., parity in TPR) by Hardt et al. [23] (POSTPRO). We employ POSTPRO because it is compatible with any base classifier, enabling fair comparisons with MISOB under identical predictive models.
Recourse methods.
We consider three widely adopted recourse methods, selected for their benchmarking popularity, available implementations, and ease of integration in an agnostic setting: Growing Spheres (GS) [31], the method by Wachter et al. [49] (WT), and the Counterfactual Conditional Heterogeneous Variational Autoencoder (CCHVAE) by Pawelczyk et al. [37]. We consider the implementations and hyperparameter settings from the benchmark by Pawelczyk et al. [38].
We assess the effectiveness of the proposed framework using overall accuracy and a range of fairness metrics, covering the whole pipeline (prediction and recourse). Specifically, we evaluate the maximum disparity across groups and the worst-group values for social burden, TPR, cost and AR. Recourse cost is quantified using the $\ell_{2}$ distance ($\times 100$) between the original instance and the recourse target. Importantly, our method operates entirely without access to sensitive attributes. Note that, to compute fairness metrics, whether for prediction or recourse, the sensitive information of instances is required, an inherent limitation of fairness metrics themselves, as discussed in Buyl and De Bie [6]. For a more detailed explanation on the experimental setting refer to Appendix B.
5.2 Results
Table 2 presents the average results over 10 random train/test splits. The findings show that training the base classifier under the MISOB framework enhances fairness both throughout the whole recourse-aware algorithmic pipeline, as well as at the prediction level, across all groups, without sacrificing overall accuracy. Notably, MISOB enhances worst-case group outcomes and, in many cases, reduces disparities across groups, achieving fairness improvements without degrading the performance for privileged groups. These results also highlight an important limitation of gap-based fairness metrics: small gaps can mask universally poor outcomes across groups, giving a misleading sense of fairness. In contrast, MISOB can produce configurations with slightly larger gaps but better metric values for every group, thereby promoting greater overall well-being. For instance, in the first block of results in Table 2, we observe that under POSTPRO, both sensitive groups achieve an equal average TPR of 0.37. In contrast, with MISOB, the average TPRs increase to 0.52 and 0.63 for the two groups, respectively. Although the gap between groups is larger, both sensitive groups achieve higher and more desirable TPR values overall.
[…]
Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR
Method strategy overall $\uparrow$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$}
race
GS - 0.81±0.02 4.56±0.01 0.03±0.02 0.27±0.03 0.08±0.08 115.69±1.19 28.37±1.95 0.06±0.01 0.06±0.02
GS PostPro 0.80±0.01 4.96±3.49 0.61±2.78 0.37±0.51 0.00±0.52 98.40±56.73 17.19±42.75 0.37±0.53 0.01±0.54
GS MISOB 0.82±0.01 3.01±0.32 0.85±0.46 0.52±0.02 0.11±0.04 93.06±0.98 27.38±1.63 0.15±0.01 0.12±0.02
WT - 0.81±0.02 1.28±0.14 0.01±0.18 0.27±0.03 0.08±0.08 38.27±0.57 12.32±0.72 0.06±0.01 0.06±0.02
WT POSTPRO 0.80±0.01 1.55±0.27 0.01±0.47 0.37±0.51 0.00±0.52 39.71±5.03 13.22±2.10 0.37±0.53 0.01±0.54
WT MISOB 0.82±0.01 0.79±0.00 0.16±0.00 0.59±0.04 0.02±0.06 30.77±0.23 11.66±0.45 0.16±0.02 0.08±0.04
CCHVAE - 0.81±0.02 6.25±0.46 0.16±0.89 0.27±0.03 0.08±0.08 119.99±3.78 27.01±7.52 0.06±0.01 0.06±0.02
CCHVAE POSTPRO 0.80±0.01 11.20±1.89 3.06±2.96 0.37±0.51 0.00±0.52 120.01±8.94 21.22±9.38 0.37±0.53 0.01±0.54
CCHVAE MISOB 0.81±0.01 4.03±0.44 0.42±0.54 0.48±0.03 0.19±0.05 105.10±1.14 27.63±1.98 0.12±0.01 0.16±0.02
gender
GS - 0.81±0.02 4.73±0.39 0.54±0.58 0.32±0.05 0.08±0.04 109.75±2.96 28.76±3.72 0.06±0.03 0.07±0.04
GS POSTPRO 0.80±0.01 24.46±25.08 13.98±8.03 0.00±0.00 0.00±0.00 145.47±47.36 28.17±63.66 0.00±0.00 0.00±0.00
GS MISOB 0.82±0.01 2.59±0.00 0.33±0.00 0.55±0.01 0.04±0.02 94.26±0.71 32.49±1.13 0.10±0.01 0.16±0.01
WT - 0.81±0.02 1.29±0.03 0.04±0.05 0.32±0.05 0.08±0.08 43.06±0.37 22.88±0.55 0.06±0.03 0.07±0.04
WT POSTPRO 0.80±0.01 1.65±0.39 0.33±0.22 0.00±0.00 0.00±0.00 42.40±3.59 20.99±2.12 0.00±0.00 0.00±0.00
WT MISOB 0.82±0.01 0.79±0.06 0.21±0.08 0.55±0.03 0.04±0.04 34.36±0.31 20.29±0.50 0.10±0.03 0.17±0.04
CCHVAE - 0.81±0.02 6.43±0.64 0.57±1.02 0.32±0.05 0.08±0.08 116.69±4.83 31.12±6.87 0.06±0.03 0.07±0.04
CCHVAE POSTPRO 0.80±0.01 12.97±4.20 6.61±4.78 0.00±0.00 0.00±0.00 119.88±7.24 27.32±12.12 0.00±0.00 0.00±0.00
CCHVAE MISOB 0.81±0.01 4.18±0.27 0.77±0.43 0.55±0.03 0.12±0.05 103.98±1.52 35.58±2.69 0.11±0.02 0.22±0.03
race & gender
GS - 0.81±0.02 5.13±0.66 1.79±1.22 0.20±0.04 0.22±0.07 124.86±2.33 48.63±3.52 0.02±0.01 0.11±0.02
GS POSTPRO 0.80±0.01 5.12±2.26 3.13±2.81 0.00±0.00 0.00±0.00 128.05±64.21 91.55±42.15 0.00±0.00 0.00±0.00
GS MISOB 0.82±0.01 3.37±0.70 1.32±1.00 0.46±0.05 0.19±0.07 109.95±2.35 51.28±3.33 0.06±0.01 0.23±0.02
WT - 0.81±0.02 1.40±0.36 0.27±0.54 0.20±0.04 0.22±0.07 49.71±1.38 30.81±2.40 0.02±0.01 0.11±0.02
WT POSTPRO 0.80±0.01 1.94±0.75 0.90±0.48 0.00±0.00 0.00±0.00 51.48±8.83 30.81±4.33 0.00±0.00 0.00±0.00
WT MISOB 0.82±0.01 0.98±0.08 0.45±0.12 0.34±0.05 0.25±0.08 41.43±0.66 28.59±0.93 0.04±0.03 0.25±0.04
CCHVAE - 0.81±0.02 7.37±0.51 3.53±1.01 0.20±0.04 0.22±0.07 133.32±1.58 53.24±3.16 0.02±0.01 0.11±0.02
CCHVAE POSTPRO 0.80±0.01 18.25±1.47 13.07±3.64 0.00±0.00 0.00±0.00 149.20±10.24 66.95±10.89 0.00±0.00 0.00±0.00
CCHVAE MISOB 0.81±0.01 5.36±0.57 2.11±1.09 0.35±0.05 0.33±0.08 116.93±0.92 52.78±1.67 0.04±0.02 0.31±0.03
Moreover, the results show that improving fairness at the prediction level does not guarantee fairer recourse outcomes and can even exacerbate them. In particular, enforcing parity in TPR through POSTPRO significantly increases the burden and cost for decision subjects. This occurs when prediction-level fairness improvements are misaligned with fairness improvements at recourse. On the contrary, the holistic perspective of MISOB enables effective improvements in group TPR, while reducing cost and burden of individuals, underscoring the importance of considering the entire pipeline. Besides, note that POSTPRO requires training a separate model for each sensitive group characterization, whereas MISOB, being blind to sensitive information, needs to be trained only once and can be evaluated across multiple group characterizations.
The effect of $C$.
The performance and fairness of MISOB are governed by the emphasis placed on high-burden individuals, controlled by the weighting parameter $C$ from Equation 4, as it can be seen in Figure 2. Increasing $C$ (i.e., the importance of high-burden instances) strengthens minimax fairness across both prediction and recourse by improving worst-group metrics: TPR and AR increase, while cost and burden decrease. However, excessively large values of $C$ can reduce overall accuracy due to the overemphasis on high-burden cases. Nonetheless, there is a broad range of $C$ values (e.g., $[0.1,0.5]$ in Figure 2(a)) where MISOB achieves substantial fairness gains with minimal impact on accuracy. Thus, $C$ governs the trade-off between fairness and accuracy, allowing users to select the value that best aligns with their objectives. In our experiments, we use $C=0.3$, which achieves strong fairness gains without compromising overall performance.
Figure 2: The performance and fairness guarantees of MISOB for different values of $C$ to compute the weighted loss in the training process for recourse methods (left) GS and (right) WT, evaluated on the Adult dataset with (a,b) race and (c,d) gender as the sensitive attributes. Increasing $C$, which enhances the emphasis on high-burden instances, strengthens minimax fairness guarantees. However, very large $C$ values may hurt overall accuracy.
6 Discussion and Limitations
We now discuss some limitations of our approach that open up further avenues for future work. Our work introduces a novel framework to compute fairness metrics in a recourse-aware setting and empirically validates the effectiveness of MISOB. Analyzing the convergence properties of the training procedure (e.g., identifying stationary points) would be a valuable direction for further research. One approach might be to frame recourse fairness as a two-player Stackelberg game, following the strategic classification perspective [8]. Additionally, applying calibration techniques [21] could improve prediction reliability (cf. Equation 3), which in turn may further enhance fairness when combined with MISOB. Finally, fairness concerns may emerge over time due to model updates, dataset shifts [7] or prediction performativity [32, 39], and recourse recommendations have been shown to be temporally fragile [12]. Extending our approach to measure and optimize the social burden of recourse in dynamic settings is, therefore, a natural path for future research. Furthermore, we aim to explore more realistic cost models that better reflect the complexities of the real world and assess how these influence fairness guarantees [46].
7 Conclusions
In this work, we study the unfairness that can arise in automated decision-making pipelines that provide recourse. We provide a theoretical characterization of the sources of unfairness in algorithmic recourse, formally linking fairness guarantees in recourse and prediction, and further highlight the limitations of the standard equal cost paradigm. To address these challenges, we introduce MISOB, a novel framework grounded in the notion of social burden, which enables a holistic treatment of fairness and goes beyond a gap-based characterization of fairness guarantees. MISOB is broadly applicable: it is compatible with any classifier–recourse pair and does not require access to sensitive attributes during training or inference. Empirical results show that improving fairness solely at the prediction level is often insufficient, and can even considerably increase the cost and burden of decision subjects. In contrast, our holistic approach leads to consistent joint fairness improvements in prediction and recourse, without compromising overall performance.
Acknowledgements
The work was partially supported by the following projects: Horizon Europe Programme, grants #101120237-ELIAS and #101120763-TANGO. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proofs of Section 3
A.1 Proof of Equation 2
Proof.
The expected recourse cost for group $\boldsymbol{s}$, computed over the entire decision pipeline (i.e., accounting for how the full population is treated throughout the process) is formally defined as:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s})}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.1)
The above expression can be decomposed into the contributions from instances receiving positive predictions and those receiving negative predictions as follows:
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot P(f(\boldsymbol{x})=0|S=\boldsymbol{s})$ (A.2)
$+\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=1)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot P(f(\boldsymbol{x})=1|S=\boldsymbol{s})$
We assume that those instances that have received a positive classification (i.e., $\boldsymbol{x}:f(\boldsymbol{x})=1$) do not change their features and, therefore, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$ for those instances. Consequently, the second term vanishes and since $P(f(\boldsymbol{x})=0|S=\boldsymbol{s})+P(f(\boldsymbol{x})=1|S=\boldsymbol{s})=1$, we get
$\mathcal{C}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot(1-P(f(\boldsymbol{x})=1|S=\boldsymbol{s}))$ (A.3)
A.2 Proof of Equation 3
Proof.
The concept of social burden analyses the effect of the algorithmic pipeline on the instances that have a true positive label:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.4)
The former expression can be decomposed as follows:
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)}$ $\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$ (A.5)
$\cdot P(f(\boldsymbol{x})=0|S=\boldsymbol{s},Y=1)$
$+\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=1)}$ $\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}$
$\cdot P(f(\boldsymbol{x})=1|S=\boldsymbol{s},Y=1)$
We assume that those individuals that have received a positive classification (i.e., $\boldsymbol{x}:f(\boldsymbol{x})=1$) do not change their features and, therefore, $\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))=0$ for those instances. Consequently, the second term vanishes and since $P(f(x)=0|S=\boldsymbol{s},Y=1)+P(f(x)=1|S=\boldsymbol{s},Y=1)=1$, we get
$\mathcal{B}_{f,g}^{\boldsymbol{s}}=$ $\mathbb{E}_{p(X|S=\boldsymbol{s},Y=1,f(\boldsymbol{x})=0)}\bigg{[}\delta((\boldsymbol{x},\boldsymbol{s}),g_{f}(\boldsymbol{x}))\bigg{]}\cdot$
$(1-P(f(\boldsymbol{x})=1|S=\boldsymbol{s},Y=1))$
Appendix B Detailed Experimental Setting
B.1 Datasets
Adult.
It is a dataset based on the data from the 1994 US Census, where the main goal is to predict whether an individual earns more than 50,000\$ per year. The 14 features used to describe the instances include occupation, marital status, and education. Furthermore, it contains sensitive information such as age, gender, and race. In our experiments, we consider sensitive groups based on gender and race, as well as intersectional groups defined by the combination of these two attributes. This dataset is publicly available in the UCI repository. [^6] http://archive.ics.uci.edu/ml/index.php Although the dataset is originally divided into predefined training and test sets, we merge them and generate new train/test splits for our empirical evaluation. We pre-process the dataset as detailed in Pawelczyk et al. [38].
Give Me Some Credit.
This dataset [^7] https://www.kaggle.com/c/GiveMeSomeCredit/ includes financial and demographic information for 150,000 individuals and is used to predict the likelihood of experiencing financial distress within two years. Each record consists of 10 feature variables, such as age, monthly income, and credit indicators, and a binary target variable indicating whether the individual experienced a serious delinquency (90+ days past due) within that timeframe (0 = Yes, 1 = No). For our fairness-aware analysis, we used age as the sensitive attribute, dividing individuals into two groups: those aged 30 or younger, and those older than 30. This distinction reflects real socioeconomic and financial differences relevant to credit risk modeling. Younger individuals often lack established credit histories and income stability, whereas older individuals tend to have more financial experience and responsibilities, which can affect their credit behavior in meaningful ways. Moreover, in our experiments, we applied the following modifications:
  1. 1.
    The original labeling assigns $Y=1$ to individuals who experienced financial distress, which is negatively framed. We flipped the labels so that $Y=1$ corresponds to a positive outcome—no financial distress within two years.
  2. 2.
    We pre-processed the dataset following the procedure of De Toni et al. [10] to ensure recourse tractability.
  3. 3.
    The original dataset is highly imbalanced, with 139,974 non-distressed cases and only 10,026 distressed ones. To avoid skewed classifiers and unreliable conclusions, we mitigate the imbalance by downsampling the majority class, resulting in an even 50/50 split between distressed and non-distressed individuals.
German Credit.
The German Credit dataset collects information about several individuals created from a German bank’s data from 1994. It contains details about the socioeconomic situation of individuals: their employment, housing, savings, etc. Besides, the set of features includes some sensitive information as well, such as gender and age. In this classification task, the objective is to predict whether an individual should obtain a good or bad credit score. This dataset is considerably smaller than the previous ones, containing only 1,000 instances and 20 features, and it is publicly available in the UCI repository. In our experiments, age has been considered as the sensitive attribute, and we have divided the instances into two (sensitive) groups based on whether their age is greater than 30 or less than or equal to 30. We pre-processed the dataset as detailed in [38].
B.2 Fairness in Prediction
We now describe the fairness-enhancing intervention selected to address fairness at the prediction level (Section 5).
Equality of opportunity by post-processing (POSTPRO).
Hardt et al. [23] propose a fairness-enhancing intervention that modifies the outcomes of a given predictor to satisfy a given fairness property defined by either equality of opportunity or equality of odds. In the case of binary predictors, their method flips the decisions with a given probability to satisfy the fairness criteria. The code for this method can be found in the AIF360888https://aif360.readthedocs.io/en/stable/ Python package.
We selected POSTPRO not only for its widespread use in fairness-in-prediction benchmarking, but also because it is compatible with any base classifier, enabling a fair comparison with MISOB.
B.3 Recourse Methods
We here describe the recourse methods used in our experimental evaluation (Section 5). For all the recourse methods, we consider the default hyperparameter settings (see Table B.1) and the implementation of the benchmark by Pawelczyk et al. [38].
Growing Spheres [31]
(GS) is a random search algorithm, which generates samples around the instance under consideration using growing hyperspheres until a point with the positive class label is found.
Differentiable Recourse [49]
(WT) proposed a method that generates recommendations by minimizing an objective function using gradient descent to find a nearby ”possible world” in which the individual receives a positive outcome, measuring proximity using the $\ell_{1}$ norm.
Counterfactual Conditional Heterogeneous Variational Autoencoder [37]
(CCHVAE) generates recommendations by ensuring that the latter maps the instance into a region of high data density that is as close as possible to the original instance, where proximity is computed in a lower-dimensional latent space defined by the VAE.
Table B.1: The hyperparameter values used in the different recourse methods.
Method Hyperparameter Type Value
GS n/a n/a n/a
WT feature_cost list None
lr float 0.01
lambda_param float 0.01
n_iter int 1000
t_max_min float 0.5
norm int 1
clamp boolean True
loss_type string ’MSE’
y_target list [0,1]
binary_cat_features boolean True
CCHVAE n_search_samples int 100
p_norm {1,2} 1
step float 0.1
max_iter int 1000
clamp boolean True
binary_cat_features boolean True
vae_params dict
- layers list [#mutable_feat, 64, 32, 8]
- train boolean True
- lambda_reg float 1e-6
- epoch int 5
- lr float 1e-3
- batch_size int 128
B.4 Additional Training Details
To mitigate overfitting during training, we employed mini-batches both in the regular training and iterative procedures for both NN and linear regression (LR). Specifically, we used a batch size of 256, a learning rate of 0.001, and the Adam optimizer. The base classifiers were trained for 6 epochs. In the case of POSTPRO, we performed the fairness-enhancing intervention after this training phase. For the MISOB framework, we implemented a two-stage training strategy: a warm-up phase of 3 epochs, followed by 3 additional epochs of iterative optimization. We experimented with longer training schedules but did not observe substantial performance gains. Therefore, we opted to limit the number of epochs to maintain training efficiency.
B.5 Additional Implementation Details
We have run all our experiments in a server with eight NVIDIA A5000 GPUs with an AMD EPYC 7252 CPU and 377GB of RAM. Nevertheless, our approach is neither RAM nor CPU/GPU intensive.
Appendix C Additional Results
In this section, we present additional experimental results using the German Credit and Give Me Some Credit datasets (see Tables C.1 and C.2, respectively). In both cases, we consider age as the sensitive attribute. Alongside NN, we also evaluate linear regression (LR) as an additional base classifier.
The results across these datasets and base classifiers reflect the same trends observed in the experiments from the main text with the Adult dataset. Specifically, MISOB consistently achieves the best worst-group performance, outperforming all other methods in terms of minimax fairness guarantees, both across the entire recourse pipeline and in the predictions of the base classifier, while incurring minimal or no loss in overall accuracy. Notably, although POSTPRO reduces TPR disparity at the base classifier level, it does so by lowering the TPR of the privileged group rather than improving that of the unprivileged group. As a result, the reduction in TPR gap comes at the cost of significantly higher burden and cost, and in some cases, reduced base accuracy.
These additional results also offer further insights. MISOB remains effective even on smaller datasets. However, we observe a higher standard deviation in performance metrics across all methods, a phenomenon also noted in the fairness literature [2], which can complicate the extraction of reliable conclusions. Finally, especially in the case of the Give Me Some Credit dataset, we note that low data quality may hinder the training of strong base classifiers. Nevertheless, MISOB is still able to significantly reduce burden and cost, even when the base classifier itself performs poorly.
[…]
Base Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR
Model method strategy overall $\uparrow$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$}
NN GS - 0.62±0.13 30.92±61.27 5.07±57.24 0.75±0.29 0.02±0.33 44.43±75.91 7.89±89.73 0.74±0.26 0.02±0.30
NN GS POSTPRO 0.61±0.03 51.46±42.93 27.59±31.75 0.05±0.09 0.01±0.12 95.31±7.06 8.73±10.26 0.04±0.07 0.02±0.12
NN GS MISOB 0.61±0.12 13.78±39.10 1.36±32.41 0.76±0.29 0.04±0.28 20.98±46.01 3.15±48.36 0.75±0.26 0.06±0.30
NN WT - 0.62±0.13 15.69±44.71 6.84±24.71 0.75±0.29 0.02±0.33 20.79±58.15 7.33±35.81 0.74±0.26 0.02±0.30
NN WT POSTPRO 0.61±0.03 27.57±31.55 2.82±29.49 0.05±0.09 0.01±0.12 48.40±36.80 3.41±24.57 0.04±0.07 0.02±0.12
NN WT MISOB 0.62±0.12 1.07±24.71 0.73±44.71 0.84±0.32 0.02±0.12 1.33±1.06 0.47±2.77 0.84±0.17 0.02±0.19
NN CCHVAE - 0.62±0.13 21.28±30.30 2.98±28.81 0.75±0.29 0.02±0.33 30.24±39.28 3.45±39.27 0.74±0.26 0.02±0.30
NN CCHVAE POSTPRO 0.61±0.03 71.10±31.93 3.72±7.92 0.05±0.09 0.01±0.12 81.10±44.04 1.51±3.07 0.04±0.07 0.02±0.12
NN CCHVAE MISOB 0.62±0.11 9.57±24.46 3.89±24.48 0.83±0.22 0.08±0.11 9.78±32.68 8.05±15.15 0.84±0.21 0.06±0.11
LR GS - 0.65±0.10 13.67±7.13 9.68±2.60 0.77±0.27 0.01±0.23 19.66±4.36 12.84±9.65 0.75±0.22 0.02±0.23
LR GS POSTPRO 0.63±0.05 32.18±24.37 3.91±16.55 0.08±0.08 0.02±0.07 73.67±42.41 10.15±27.05 0.05±0.02 0.02±0.07
LR GS MISOB 0.63±0.06 5.09±5.16 3.97±1.12 0.80±0.27 0.05±0.23 7.69±4.36 5.28±9.65 0.82±0.17 0.02±0.09
LR WT - 0.65±0.10 1.14±2.39 0.45±1.21 0.77±0.27 0.01±0.23 1.70±2.72 0.64±1.81 0.75±0.22 0.02±0.23
LR WT POSTPRO 0.63±0.05 5.83±6.90 3.58±2.84 0.08±0.08 0.02±0.07 10.15±12.59 7.25±3.59 0.05±0.02 0.02±0.07
LR WT MISOB 0.67±0.05 0.11±0.32 0.11±0.32 0.92±0.12 0.01±0.11 0.11±0.32 0.11±0.32 0.91±0.10 0.02±0.11
LR CCHVAE - 0.65±0.10 15.68±22.58 4.30±19.30 0.77±0.27 0.01±0.23 27.28±39.48 4.03±32.76 0.75±0.22 0.02±0.23
LR CCHVAE POSTPRO 0.63±0.05 56.63±26.03 3.31±35.21 0.08±0.08 0.02±0.07 96.10±56.92 5.58±19.81 0.05±0.02 0.02±0.07
LR CCHVAE MISOB 0.66±0.09 4.50±16.61 3.02±2.81 0.84±0.23 0.05±0.28 12.52±34.00 8.87±28.04 0.86±22.67 0.05±16.57
[…]
Base Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR
Model Method strategy overall $\uparrow$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$} worst $\downarrow$} $\Delta$ $|\downarrow|$} worst $\uparrow$} $\Delta$ $|\downarrow|$}
NN GS - 0.56±0.02 3.27±0.88 0.20±0.56 0.30±0.11 0.33±0.12 25.04±7.29 3.79±2.37 0.24±0.09 0.26±0.10
NN GS POSTPRO 0.53±0.01 5.32±3.08 3.79±1.11 0.25±0.18 0.01±0.04 23.40±7.51 2.89±7.62 0.32±0.40 0.13±0.10
NN GS MISOB 0.56±0.03 1.65±1.82 0.49±1.00 0.38±0.24 0.37±0.27 10.30±6.17 0.01±4.44 0.33±0.22 0.34±0.23
NN WT - 0.56±0.02 1.78±0.71 0.50±0.51 0.30±0.11 0.33±0.12 14.39±2.49 6.24±1.92 0.24±0.09 0.26±0.10
NN WT POSTPRO 0.53±0.01 2.55±2.17 0.84±4.97 0.25±0.18 0.01±0.04 13.62±4.73 5.79±7.06 0.32±0.40 0.13±0.10
NN WT MISOB 0.59±0.05 1.32±1.64 0.31±1.19 0.55±0.13 0.24±0.11 11.37±6.71 4.47±4.25 0.46±0.10 0.20±0.16
NN CCHVAE - 0.56±0.02 6.43±3.48 1.04±3.43 0.30±0.11 0.33±0.12 35.00±13.03 1.49±17.43 0.24±0.09 0.26±0.10
NN CCHVAE POSTPRO 0.53±0.01 20.62±24.14 10.17±17.64 0.25±0.18 0.01±0.04 29.21±35.16 7.57±31.21 0.32±0.40 0.13±0.10
NN CCHVAE MISOB 0.57±0.04 4.15±2.40 0.76±2.10 0.43±0.16 0.31±0.13 24.25±11.47 1.19±9.52 0.36±0.15 0.28±0.14
LR GS - 0.51±0.09 10.95±12.15 5.91±1.98 0.17±0.25 0.03±0.05 33.38±26.30 13.63±8.47 0.16±0.18 0.03±0.03
LR GS POSTPRO 0.54±0.10 17.31±11.86 6.21±2.30 0.11±0.23 0.01±0.03 39.23±30.92 6.55±6.33 0.09±0.18 0.04±0.02
LR GS MISOB 0.52±0.02 2.59±1.57 1.51±1.32 0.45±0.24 0.04±0.15 16.61±0.04 4.60±1.64 0.34±0.36 0.06±0.17
LR WT - 0.51±0.09 2.75±2.12 0.77±0.87 0.17±0.25 0.03±0.05 10.63±6.29 4.74±1.63 0.16±0.18 0.03±0.03
LR WT POSTPRO 0.54±0.10 5.25±1.43 0.89±0.76 0.11±0.23 0.01±0.03 16.16±4.44 7.93±2.48 0.09±0.18 0.04±0.02
LR WT MISOB 0.56±0.11 1.63±0.30 0.90±0.57 0.54±0.28 0.04±0.35 6.67±6.28 3.11±1.59 0.51±0.27 0.01±0.36
LR CCHVAE - 0.51±0.09 14.81±11.34 3.26±10.71 0.17±0.25 0.03±0.05 48.08±17.34 7.77±35.51 0.16±0.18 0.03±0.03
LR CCHVAE POSTPRO 0.54±0.10 17.51±25.99 6.30±16.01 0.11±0.23 0.01±0.03 56.87±39.04 3.33±28.95 0.09±0.18 0.04±0.02
LR CCHVAE MISOB 0.54±0.07 7.55±5.73 5.50±10.98 0.44±0.38 0.01±0.39 29.61±34.15 8.61±37.65 0.44±0.36 0.04±0.37